Room temperature constant 20 ° C and relative humidity 20%. How many grams of water must be added to the room air to obtain relative humidity to rise to 60%?
1
Expert's answer
2019-03-11T12:30:58-0400

Solution:

Relative humidity is the ratio of vapor density to saturation vapor density.

In the first case:


"\phi_1=\frac{\rho_1} {\rho_0} 100\%"

Vapor density equals to mass of vapor divided by the volume of the room:


"\rho_1=\frac{m_1} V"

Then:

"\phi_1=\frac{m_1} {\rho_0 V} 100\%"

Therefore:

"m_1=\frac{\phi_1\rho_0 V} {100\%}"

Similarly we can find mass of vapor after adding the water:


"\phi_2=\frac{\rho_2} {\rho_0} 100\%=\frac{m_2} {\rho_0 V} 100\%"

"m_2=\frac{\phi_2\rho_0 V} {100\%}"

Mass of water that must be added to the room air:


"\Delta{m}=\frac{\rho_0 V} {100\%} (\phi_2-\phi_1)"

Saturation vapor density at "20^\circ\:C:"


"\rho_0=17.3\:g/m^3"

Then:

"\Delta{m}=\frac{17.3 V} {100\%} (60\%-20\%)=6.92V\:(g)"

Answer:

"\Delta{m}=6.92V\:g"